the lenght of a rectangle is 4 inches greater thatn the width the perimeter of the recatangle is 24 inches find the dimensions of the rectangle
Answer by doukungfoo(195)

Perimeter of a rectangle is 2 times the length plus 2 times the width
written as a formula it looks like this
P = 2L + 2W
Where P is perimeter
L is length and
W is width
Ok so back to the rectangle in this problem
Lets call the width x
ok so
width = x
Now we are told that the length is 4 inches greater than the width
ok well if the width is x then
the length is x + 4
Finally we are told that the perimeter is 24 inches
so lets put all this into the perimeter formula
P = 2L + 2W
24 = 2(x+4) + 2(x)
24 = 2x + 8 + 2x
24 = 4x + 8
16 = 4x
4 = x
Answer: the width is 4 inches
now use this to find the length
length is x + 4
so
4 + 4 = 8
Answer: the length is 8 inches
Check both answers in the perimeter formula
P = 2L + 2W
24 = 2(8) + 2(4)
24 = 16 + 8
24 = 24